# A framework for fluctuating times and counting observables in stochastic excursions

### v1

@@ III.1 Characteristic function
[…]
$=C\langle y_{A}|W_{AB\bm{\xi}}(-\mathbb{W}_{\!B\bm{\xi}}-i\omega)^{-1}W_{BA\bm{\xi}}|x_{A}\rangle.$
@@
$+i\omega E(\hat{\mathcal{T}})-\frac{\omega^{2}}{2}E(\hat{\mathcal{T}}^{2})-\sum_{\alpha=1}^{r}\omega\xi_{\alpha}E(\hat{\mathcal{T}}\hat{Q}_{\alpha})+\ldots.$
@@
$G(\bm{\xi},\omega)=$ $\ C\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|x_{A}\rangle+i\omega C\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{2}W_{BA\bm{\xi}}|x_{A}\rangle$ (24)
$-\omega^{2}C\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{3}W_{BA\bm{\xi}}|x_{A}\rangle+\ldots.$
@@ III.2 Moments of the excursion time
The first two moments of the duration of excursions $\hat{\mathcal{T}}$ are obtained by setting $\bm{\xi}=0$ in Eq. (24) and comparing with Eq. (23):
$E(\hat{\mathcal{T}})$ $=C\langle y_{A}|W_{AB}M^{2}W_{BA}|x_{A}\rangle,$ (25)
$E(\hat{\mathcal{T}}^{2})$ $=2C\langle y_{A}|W_{AB}M^{3}W_{BA}|x_{A}\rangle.$ (26)
@@
$E(\hat{\mathcal{T}}^{n})$ $=n!C\langle y_{A}|W_{AB}M^{n+1}W_{BA}|x_{A}\rangle$ (28)
@@ III.3 Moments of the counting observables
For general counting observables, we first expand the tilted matrices (18) over $\bm{\xi}$:
$\mathbb{W}_{\!\bm{\xi}}=\mathbb{W}+i\sum_{\alpha=1}^{r}\xi_{\alpha}W^{(\alpha)}-\frac{1}{2}\sum_{\alpha,\beta}^{r}\xi_{\alpha}\xi_{\beta}W^{(\alpha,\beta)},$ (29)
@@
Notice these are element-wise (Hadamard) products, not matrix products. Moreover, this is defined for the transition matrices $W$ which have zeros in the diagonal, and not the full matrix $\mathbb{W}$ in Eq. (3). Carrying out a similar expansion for $M_{\bm{\xi}}$:
$M_{\bm{\xi}}=$ $\ M+i\sum_{\alpha=1}^{r}\xi_{\alpha}MW_{B}^{(\alpha)}M-\frac{1}{2}\sum_{\alpha,\beta=1}^{r}\xi_{\alpha}\xi_{\beta}\bigg{(}MW_{B}^{(\alpha,\beta)}M$ (32)
@@
$E(\hat{Q}_{\alpha})$ $=C\leavevmode\nobreak\ \bra{y_{A}}\Big{(}W_{AB}^{(\alpha)}MW_{BA}+W_{AB}MW_{BA}^{(\alpha)}+W_{AB}MW_{B}^{(\alpha)}MW_{BA}\Big{)}\ket{x_{A}},$ (33)
$E(\hat{Q}_{\alpha}\hat{Q}_{\beta})$ $=C\leavevmode\nobreak\ \langle y_{A}|\Bigg{\{}W_{AB}^{(\alpha,\beta)}MW_{BA}+W_{AB}MW_{BA}^{(\alpha,\beta)}+W_{AB}MW_{B}^{(\alpha,\beta)}MW_{BA}$ (34)
@@
The interpretation of Eq. (33) is intuitive: the first and second terms are the average of $\hat{Q}_{\alpha}$ in the transitions from $A\to B$ and then from $B\to A$. The last term is then the average of $\hat{Q}_{\alpha}$ while inside $B$. The interpretation of Eq. (34) is similar albeit more technical. From Eqs. (33) and (34) one can also readily build the covariance matrix
@@ III.4 Correlation between excursion duration and counting observables
Very often the counting observables and the excursion duration are correlated. This is clearly seen in the expansion on Eqs. (23) and (24). To calculate $E(\hat{\mathcal{T}}\hat{Q}_{\alpha})$, we perform another series expansion $i\omega C\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{2}W_{BA\bm{\xi}}|x_{A}\rangle$ to first order in $\bm{\xi}$. This results in
$E(\hat{\mathcal{T}}\hat{Q}_{\alpha})=$ $\ C\leavevmode\nobreak\ \langle y_{A}|\Big{\{}W_{AB}^{(\alpha)}M^{2}W_{BA}+W_{AB}M^{2}W_{BA}^{(\alpha)}$ (36)
@@
From this result and the averages of excursion duration [Eq. (25)] and generic counting observables [Eq. (33)], one can quantitatively explore the covariance between a generic counting observable and the excursion duration.
@@ III.5 Conditional moments
At last, another important calculation that arises in the context of excursions is when a counting observable is conditioned. For simplicity, let us consider a single counting observable $\hat{Q}_{\alpha}$. We first calculate $E(\hat{Q}_{\alpha}|\hat{\mathcal{T}}=t)$. Taking the Fourier transform of Eq. (19) with respect only to $\xi$ yields the conditional characteristic function
@@
The first order term in a series expansion in $\xi$ will give us $i\xi E(\hat{Q}_{\alpha}|\hat{\mathcal{T}}=t)$. For $W_{AB\xi}$ and $W_{BA\xi}$ we use Eq. (29). For the exponential we use the Baker-Campbell-Hausdorff formula
@@
Putting all the above equations together, we find
$E(\hat{Q}_{\alpha}|\hat{\mathcal{T}}=t)=$ $\ \frac{C}{P(t)}\Bigg{\{}\langle y_{A}|W_{AB}^{(\alpha)}e^{\mathbb{W}_{B}t}W_{BA}|x_{A}\rangle$ (39)
$+\langle y_{A}|W_{AB}e^{\mathbb{W}_{B}t}W_{BA}^{(\alpha)}|x_{A}\rangle$
$+\int\limits_{0}^{t}dt^{\prime}\leavevmode\nobreak\ \langle y_{A}|W_{AB}e^{\mathbb{W}_{B}(t-t^{\prime})}W_{B}^{(\alpha)}e^{\mathbb{W}_{B}t^{\prime}}W_{BA}|x_{A}\rangle\Bigg{\}}.$
This quantity is useful for inferring the value of the counting observable from the excursion time. The reason is because the conditional mean is exactly the optimal mean-squared predictor; i.e., the function of $\theta(\hat{T})$ that minimizes the mean-squared error between $E\Big{[}\big{(}\hat{Q}_{\alpha}-\theta(\hat{T})\big{)}^{2}\Big{]}$.
@@
$\theta_{\rm lin}(\hat{\mathcal{T}})=E(\hat{Q}_{\alpha})+\frac{{\rm cov}(\hat{Q}_{\alpha},\hat{\mathcal{T}})}{{\rm var}(\hat{\mathcal{T}})}\Big{(}\hat{\mathcal{T}}-E(\hat{\mathcal{T}})\Big{)}.$ (40)
This quantity can be straightforwardly calculated from Eqs. (25), (27), (33) and (36). This result is a type of inference that can be obtained from the excursion framework by assessing the statistics of excursion times. Inference of counting observables such as entropy production in hidden/coarse grained settings is a quite active area rahavFluctuationRelationsCoarsegraining2007 ; boEntropyProductionStochastic2014 ; biskerHierarchicalBoundsEntropy2017 ; yuDissipationLimitedResolutions2024a ; harunariWhatLearnFew2022 ; espositoStochasticThermodynamicsCoarse2012 ; ehrichTightestBoundHidden2021 ; blomMilestoningEstimatorsDissipation2024 ; liangThermodynamicBoundsSymmetry2024 ; Tan2021 , it would be interesting to further explore what insights Eq. (40) can bring in this context.
@@ III.6 Connection with full counting statistics steady state
We now discuss how to use the results developed in this section to connect our framework with full counting statistics (FCS). Let $\hat{\mathcal{Q}}$ denote a FCS counting observable. This counting observable is defined in a similar way to Eq. (15), but considered over a fixed time interval $[0,t]$ rather than per excursion. In the limit of long time $t$, one then defines the average current and diffusion coefficient (also called noise):
@@
Let us assume region $A$ is spanned by a single state, this means that excursions become statistically independent of each other. The connection between FCS results and counting observables at the excursion level comes from a concatenation of several excursions, all starting in the same state $x_{A}$. Let $\hat{Q}_{n},\leavevmode\nobreak\ \hat{\mathcal{T}}_{n}$ denote a counting observable and the excursion time during the $n$-th excursion. The total cycle time of an excursion also includes the residence time $\hat{\tau}_{x,n}$ (i.e. the time the system spends in $x$ after the $(n-1)$-th excursion is over but before the $n$-th excursion has started), so it is defined as $\hat{\mathcal{T}}_{n}^{\rm cyc}=\hat{\mathcal{T}}_{n}+\hat{\tau}_{x,n}$. Note that $\hat{\tau}_{x,n}$ and $\hat{\mathcal{T}}_{n}$ are statistically independent. Finally, let $\hat{\mathcal{N}}(t)$ denote the number of excursions that took place in the interval $[0,t]$. This random variable forms a renewal process cox1967renewal . For sufficiently large $t$, we can then write
@@
In Ref. Fiusa2025 we showed that the average current and noise from Eq. (41) could be rewritten as
@@
$D$ $=\frac{{\rm var}(\hat{Q})}{\mu}+\frac{\Delta^{2}}{\mu^{3}}E(\hat{Q})^{2}-\frac{2E(\hat{Q})}{\mu^{2}}{\rm cov}(\hat{Q},\hat{\mathcal{T}}),$ (45)
where we defined $\mu=E(\hat{\mathcal{T}}_{n}^{\rm cyc})=E(\hat{\mathcal{T}})+\Gamma_{x}^{-1}$ and $\Delta^{2}={\rm var}(\hat{\mathcal{T}}_{n}^{\rm cyc})={\rm var}(\hat{\mathcal{T}})+\Gamma_{x}^{-2}$ as the mean and variance of $\hat{\mathcal{T}}^{\rm cyc}$. Residence times $\hat{\tau}_{x}$ are taken to be exponentially distributed with parameter $\Gamma_{x}$, and since they are statistically independent from $\hat{Q}$ and $\hat{\mathcal{T}}$, it follows that ${\rm cov}(\hat{Q},\hat{\mathcal{T}}^{\rm cyc})={\rm cov}(\hat{Q},\hat{\mathcal{T}})$.
The results in Eqs. (44) and (45) are one of the main achievements of the excursion framework. It provides a decomposition of the FCS noise using only statistics at the level of excursions. With the tools developed in this section, such as Eqs. (25), (27), (33), (34), and (36), the average current and the noise can be readily computed.
@@ IV.1 Dynamical activity
$G(\xi)=C\sum_{n=0}^{\infty}\langle y_{A}|W_{AB}\Gamma_{B}^{-1}\psi^{n}W_{BA}|x_{A}\rangle e^{i(n+2)\xi}.$ (50)
@@
$P\big{(}\hat{\mathcal{A}}=n+2\big{)}=C\leavevmode\nobreak\ \langle y_{A}|W_{AB}\Gamma_{B}^{-1}\psi^{n}W_{BA}|x_{A}\rangle.$ (51)
@@
$E\big{(}\hat{\mathcal{A}}\big{)}$ $=\sum_{n=0}^{\infty}(n+2)P\big{(}\hat{\mathcal{A}}=n+2\big{)}$ (52)
$=2+C\langle y_{A}|W_{AB}\Gamma_{B}^{-1}(1-\psi)^{-2}\psi W_{BA}|x_{A}\rangle.$
@@
where $E(\hat{\mathcal{A}}_{B}):=C\langle y_{B}|W_{AB}MW_{\!B}MW_{BA}|x_{A}\rangle$ is the average number of transitions within region $B$ only. A similar calculation yields the variance of the number of jumps
${\rm var}\big{(}\hat{\mathcal{A}}\big{)}=$ $\ 2C\langle y_{B}|W_{AB}MW_{B}MW_{\!B}MW_{BA}|x_{A}\rangle$

### v2

@@ III.1 Characteristic function
$G(\bm{\xi},\omega)$ $=\int dq_{1}\ldots dq_{r}\int\limits_{0}^{\infty}dt\penalty 10000\ e^{-i(\bm{q}\cdot\bm{\xi}+\omega t)}P(\bm{q},t)$ (22)
$=C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB\bm{\xi}}(-\mathbb{W}_{\!B\bm{\xi}}-i\omega)^{-1}W_{BA\bm{\xi}}|x_{A}\rangle.$
@@
$+i\omega E(\hat{T})-\frac{\omega^{2}}{2}E(\hat{T}^{2})-\sum_{\alpha=1}^{r}\omega\xi_{\alpha}E(\hat{T}\hat{Q}_{\alpha})+\ldots.$
@@
$G(\bm{\xi},\omega)=\ C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}W_{BA\bm{\xi}}|x_{A}\rangle+i\omega C_{x_{A}\to y_{A}}$ (24)
$\times\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{2}W_{BA\bm{\xi}}|x_{A}\rangle-\omega^{2}C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{3}W_{BA\bm{\xi}}|x_{A}\rangle+\ldots.$
@@ III.2 Moments of the excursion time
The first two moments of the duration of excursions $\hat{T}$ are obtained by setting $\bm{\xi}=0$ in Eq. (24) and comparing with Eq. (23):
[…]
$E(\hat{T}^{2})$ $=2C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}M^{3}W_{BA}|x_{A}\rangle.$ (26)
@@
Alternatively, any moment of the excursion time can also be easily computed directly from Eq. (14) and the identity $\int_{0}^{\infty}dt\penalty 10000\ t^{n}e^{\mathbb{W}_{\!B}t}=(-1)^{n+1}n!\mathbb{W}_{\!B}^{-(n+1)}$, leading to
$E(\hat{T}^{n})$ $=n!C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}M^{n+1}W_{BA}|x_{A}\rangle$ (28)
@@ III.3 Moments of the counting observables
For general counting observables, we first expand the tilted matrices [c.f. (18)] over $\bm{\xi}$:
$\mathbb{W}_{\!\bm{\xi}}\simeq\mathbb{W}+i\sum_{\alpha=1}^{r}\xi_{\alpha}W^{(\alpha)}-\frac{1}{2}\sum_{\alpha,\beta}^{r}\xi_{\alpha}\xi_{\beta}W^{(\alpha,\beta)},$ (29)
@@
Notice these are element-wise (Hadamard) products, not matrix products. We keep the expansion up to order two since we are only interested in the average and variance. To compute higher order momenta, such as skewness or kurtosis, all one has to do is to keep higher order terms. Moreover, this is defined for the transition matrices $W$ which have zeros in the diagonal, and not the full matrix $\mathbb{W}$ in Eq. (3). Carrying out a similar expansion for $M_{\bm{\xi}}$:
$M_{\bm{\xi}}\simeq$ $\ M+i\sum_{\alpha=1}^{r}\xi_{\alpha}MW_{B}^{(\alpha)}M-\frac{1}{2}\sum_{\alpha,\beta=1}^{r}\xi_{\alpha}\xi_{\beta}\bigg(MW_{B}^{(\alpha,\beta)}M$ (32)
@@
[…]
$E(\hat{Q}_{\alpha}\hat{Q}_{\beta})$ $=C_{x_{A}\to y_{A}}\penalty 10000\ \langle y_{A}|\Bigg\{W_{AB}^{(\alpha,\beta)}MW_{BA}+W_{AB}MW_{BA}^{(\alpha,\beta)}+W_{AB}MW_{B}^{(\alpha,\beta)}MW_{BA}$ (34)
@@
The interpretation of Eq. (33) is intuitive: the first and second terms are the average of $\hat{Q}_{\alpha}$ in the transitions from $A\to B$ and then from $B\to A$. The last term is then the average of $\hat{Q}_{\alpha}$ while inside $B$. The interpretation of Eq. (34) is similar albeit more technical. Note that the results (33) and (34) are exact. The expansion over the counting fields is kept up to second order in Eq. (32) because higher order terms do not contribute to the first and second momenta. The variance is readily evaluated by
${\rm var}(\hat{Q}_{\alpha})=E(\hat{Q}_{\alpha}^{2})-E(\hat{Q}_{\alpha})^{2},$ (35)
where $E(\hat{Q}_{\alpha}^{2})$ is given by Eq. (34) with $\beta=\alpha$. From Eqs. (33) and (34) one can also readily build the covariance matrix
@@ III.4 Correlation between excursion duration and counting observables
Very often the counting observables and the excursion duration are correlated. This is clearly seen in the expansion on Eqs. (23) and (24). To calculate $E(\hat{T}\hat{Q}_{\alpha})$, we perform another series expansion $i\omega C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB\bm{\xi}}M_{\bm{\xi}}^{2}W_{BA\bm{\xi}}|x_{A}\rangle$ to first order in $\bm{\xi}$. This results in
$E(\hat{T}\hat{Q}_{\alpha})=$ $\ C_{x_{A}\to y_{A}}\penalty 10000\ \langle y_{A}|\Big\{W_{AB}^{(\alpha)}M^{2}W_{BA}+W_{AB}M^{2}W_{BA}^{(\alpha)}$ (37)
@@
From this result and the averages of excursion duration [Eq. (25)] and generic counting observables [Eq. (33)], one can quantitatively explore the covariance between a generic counting observable and the excursion duration. Moreover, results in this section provide a systematic way to evaluate the average and fluctuations of any counting observable in an excursion as well as excursion durations with simple closed-formula expressions that depend only on elements of the stochastic transition matrix (3).
@@ III.5 Conditional moments
At last, another important calculation that arises in the context of excursions is when a counting observable is conditioned. For simplicity, let us consider a single counting observable $\hat{Q}_{\alpha}$. We first calculate $E(\hat{Q}_{\alpha}|\hat{T}=t)$. Taking the Fourier transform of Eq. (19) with respect only to $\xi$ yields the conditional characteristic function
@@
The first order term in a series expansion in $\xi$ will give us $i\xi E(\hat{Q}_{\alpha}|\hat{T}=t)$. For $W_{AB\xi}$ and $W_{BA\xi}$ we use Eq. (29). For the exponential we use the Baker-Campbell-Hausdorff formula
$e^{\mathbb{W}_{B\xi}t}\simeq e^{\mathbb{W}_{B}t}+i\xi\int\limits_{0}^{t}dt^{\prime}\penalty 10000\ e^{\mathbb{W}_{B}(t-t^{\prime})}W_{B}^{(\alpha)}e^{\mathbb{W}_{B}t^{\prime}}.$ (39)
Putting all the above equations together, we find
$E(\hat{Q}_{\alpha}|\hat{T}=t)=$ $\ \frac{C_{x_{A}\to y_{A}}}{P(t)}\Bigg\{\langle y_{A}|W_{AB}^{(\alpha)}e^{\mathbb{W}_{B}t}W_{BA}|x_{A}\rangle$ (40)
$+\langle y_{A}|W_{AB}e^{\mathbb{W}_{B}t}W_{BA}^{(\alpha)}|x_{A}\rangle$
$+\int\limits_{0}^{t}dt^{\prime}\penalty 10000\ \langle y_{A}|W_{AB}e^{\mathbb{W}_{B}(t-t^{\prime})}W_{B}^{(\alpha)}e^{\mathbb{W}_{B}t^{\prime}}W_{BA}|x_{A}\rangle\Bigg\}.$
@@
$).\theta_{\rm lin}(\hat{T})=E(\hat{Q}_{\alpha})+\frac{{\rm cov}(\hat{Q}_{\alpha},\hat{T})}{{\rm var}(\hat{T})}\Big(\hat{T}-E(\hat{T}$$\Big)$. (41)
This quantity can be straightforwardly calculated from Eqs. (25), (27), (33) and (37). This result is a type of inference that can be obtained from the excursion framework by assessing the statistics of excursion times. Inference of counting observables such as entropy production in hidden/coarse grained settings is a quite active area rahavFluctuationRelationsCoarsegraining2007 ; boEntropyProductionStochastic2014 ; biskerHierarchicalBoundsEntropy2017 ; yuDissipationLimitedResolutions2024a ; harunariWhatLearnFew2022 ; espositoStochasticThermodynamicsCoarse2012 ; ehrichTightestBoundHidden2021 ; blomMilestoningEstimatorsDissipation2024 ; liangThermodynamicBoundsSymmetry2024 ; Tan2021 , it would be interesting to further explore what insights Eq. (41) can bring in this context, especially because inference of counting observables beyond entropy production is a very difficult task van_der_Meer2022 ; van_der_Meer2023 .
@@ III.6 Connection with full counting statistics steady state
We now discuss how to use the results developed in this section to connect our framework with full counting statistics (FCS), which is a framework developed to address current fluctuations and statistics of integrated charges schaller2014open ; landi2024a ; Campisi2011 ; Esposito2007 ; Esposito2009 ; Brandes2008 ; Saito2008 . Let $\hat{\mathcal{Q}}$ denote a FCS counting observable. This counting observable is defined in a similar way to Eq. (15), but considered over a fixed time window $[0,t]$. That is,
$\hat{\mathcal{Q}}(t)=\sum_{\ell xy}\nu_{\ell xy}\hat{N}_{\ell xy}(t),$ (42)
where the random variable $\hat{N}_{\ell xy}(t)$ counts how many times the transition $y\to_{\ell}x$ took place within the time time window. In the limit of long time $t$, the average current and the diffusion coefficient (also called noise) of this observable are defined as
@@
We will show next how to recover the same results but using counting observables defined per excursion and without assessing the long time limit steady state.
Let us assume region $A$ is spanned by a single state, this means that excursions become statistically independent of each other. The connection between FCS results and counting observables at the excursion level comes from a concatenation of several excursions, all starting in the same state $x_{A}$. Let $\hat{Q}_{n},\penalty 10000\ \hat{T}_{n}$ denote a counting observable and the excursion time during the $n$-th excursion. The total cycle time of an excursion also includes the residence time $\hat{\tau}_{x,n}$ (i.e. the time the system spends in $x$ after the (n1)(n-1)-th excursion is over but before the $n$-th excursion has started), so it is defined as $\hat{T}_{n}^{\rm cyc}=\hat{T}_{n}+\hat{\tau}_{x,n}$. Note that $\hat{\tau}_{x,n}$ and $\hat{T}_{n}$ are statistically independent. Finally, let $\hat{\mathcal{N}}(t)$ denote the number of excursions that took place in the interval $[0,t]$. This random variable forms a renewal process cox1967renewal . For sufficiently large $t$, we can then write
@@
In Ref. Fiusa2025 we showed that the average current and noise from Eq. (43) can be obtained as
@@
[…]
The results in Eqs. (46) and (47) are one of the main achievements of the excursion framework. First, it provides a way to compute steady state quantities in FCS by looking at statistics of individual excursions, which are subtrajectories and therefore easier to treat. Second, it unravels a decomposition of the FCS noise into three distinct components, shedding light into the mechanisms behind noise generation in classical stochastic dynamics. Investigating this further, and especially under the lenses of uncertainty relations, can provide fundamental insights into the nature of current precision. With the tools developed in this section, such as Eqs. (25), (27), (33), (34), (35), and (37), the average current and the noise can be readily computed.
@@ IV.1 Dynamical activity
$G(\xi)=C_{x_{A}\to y_{A}}\sum_{n=0}^{\infty}\langle y_{A}|W_{AB}\Gamma_{B}^{-1}\psi^{n}W_{BA}|x_{A}\rangle e^{i(n+2)\xi}.$ (52)
@@
$P\big(\hat{\mathcal{A}}=n+2\big)=C_{x_{A}\to y_{A}}\penalty 10000\ \langle y_{A}|W_{AB}\Gamma_{B}^{-1}\psi^{n}W_{BA}|x_{A}\rangle.$ (53)
@@
[…]
$=2+C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}\Gamma_{B}^{-1}(1-\psi)^{-2}\psi W_{BA}|x_{A}\rangle.$
@@
where $E(\hat{\mathcal{A}}_{B}):=C_{x_{A}\to y_{A}}\langle y_{A}|W_{AB}MW_{\!B}MW_{BA}|x_{A}\rangle$ is the average number of transitions within region $B$ only. A similar calculation yields the variance of the number of jumps
[…]